40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 28 02:33:04 2011, Jenne, Update, Locking, Yarm flashing, but not yet locked 
    Reply  Thu Jul 28 16:10:34 2011, kiwamu, Update, LSC, Both arm locked 
Message ID: 5051     Entry time: Thu Jul 28 02:33:04 2011     Reply to this: 5054
Author: Jenne 
Type: Update 
Category: Locking 
Subject: Yarm flashing, but not yet locked 

Because I'm too lazy to write a cohenrent elog right now, here's my notes that I wrote while working tonight:

Elog notes, 27July2011

Aligned Xarm, just to check on it.  Had to flip sign of TRX in DCPD filter bank (to gain of -1) to make the signal positive.

Restored Yarm, see some slight flashing, but no lock yet.
Adjusted phase rotation of AS55 from 56.5deg to 60deg, just by-eye trying to maximize AS55I, my arm error signal. AS55I goes from ~ -40 to +60 counts

Tried fitzing with Yarm gain, flipping sign, incr gain. No real change in signals, or flashing.


Incr. ETMY oplev gains to -0.4 from -0.2
Engaged ELP35's on Pit and Yaw, to be more similar to other optics.  However, right now all of the optics have different things in their filter banks.  Why??

Arm is flashing pretty reliably now, but still not locking.  The trigger threshold is always satisfied, so that's not it.

ELOG V3.1.3-